# Towards Trustworthy Reranking: A Simple yet Effective Abstention Mechanism

### v4

@@ Abstract
Neural Information Retrieval (NIR) has significantly improved upon heuristic-based IR systems. Yet, failures remain frequent, the models used often being unable to retrieve documents relevant to the user’s query. We address this challenge by proposing a lightweight abstention mechanism tailored for real-world constraints, with particular emphasis placed on the reranking phase. We introduce a protocol for evaluating abstention strategies in a black-box scenario, demonstrating their efficacy, and propose a simple yet effective data-driven mechanism. We provide open-source code for experiment replication and abstention implementation, fostering wider adoption and application in diverse contexts.
@@ 1 Introduction
In recent years, Neural Information Retrieval (NIR) has emerged as a promising approach to addressing the challenges of Information Retrieval (IR) on various tasks Guo et al. (2016); Mitra and Craswell (2017); Zhao et al. (2022). Central to the NIR paradigm are the pivotal stages of retrieval and reranking Robertson et al. (1995); Ni et al. (2021); Neelakantan et al. (2022), which collectively play a fundamental role in shaping the performance and outcomes of IR systems Thakur et al. (2021). While the objective of the retrieval stage is to efficiently fetch candidate documents from a vast corpus based on a user’s query (frequentist Ramos et al. (2003); Robertson et al. (2009) or bi-encoder-based approaches Karpukhin et al. (2020)), reranking aims at reordering these retrieved documents in a slower albeit more effective way, using more sophisticated techniques (bi-encoder- or cross-encoder-based Nogueira and Cho (2019); Khattab and Zaharia (2020)) [^1] The most commonly used approaches are to use a frequentist method followed by a bi-encoder (light option) or a bi-encoder followed by a cross-encoder (heavier option)..
@@
However, despite advancements in NIR techniques, retrieval and reranking processes frequently fail, due for instance to the quality of the models being used, to ambiguous user queries or insufficient relevant documents in the database. These can have unsuitable consequences, particularly in contexts like Retrieval-Augmented Generation (RAG) Lewis et al. (2020), where retrieved information is directly used in downstream tasks, thus undermining the overall reliability of the system Yoran et al. (2023); Wang et al. (2023); Baek et al. (2023).
@@
Recognizing the critical need to address these challenges, the implementation of abstention mechanisms emerges as a significant promise within the IR landscape. This technique offers a pragmatic approach by abstaining from delivering results when the model shows signs of uncertainty, thus ensuring users are not misled by potentially erroneous information. Although the subject of abstention in machine learning has long been of interest Chow (1957); El-Yaniv and Wiener (2010); Geifman and El-Yaniv (2017, 2019), most research has focused on the classification setting, leaving lack of significant initiatives in IR. The few related work proposes computationally intensive learning strategies Cheng et al. (2010); Mao et al. (2023), making them difficult to implement in the case of NIR, where the models can have up to billions of parameters.
@@
Contributions. In this paper, we explore lightweight abstention mechanisms that adhere to realistic industrial constraints encountered in practice, namely (i) black-box access-only to query-documents relevance scores (compatible with API services), (ii) marginal computational and latency overhead, (iii) configurable abstention rates depending on the application. In line with industrial applications of IR systems for RAG or search engines, our method focuses on assessing the reranking quality of the top retrieved documents.
Our key contributions are:
@@
1. We devise a detailed protocol for evaluating abstention strategies in a black-box reranking setting and demonstrate their relevance.
@@
2. We devise a simple yet effective reference-based abstention mechanism, outperforming reference-free heuristics at zero incurred cost. We perform ablations and analyses on this method and are confident about its potential in practical settings.
@@
3. We release a code package [^2] https://github.com/artefactory/abstention-reranker.git, under MIT license. and artifacts [^3] https://huggingface.co/datasets/manu/reranker-scores. to enable full replication of our experiments and the implementation of plug-and-play abstention mechanisms for any use cases.
@@ 2.1 Notations
General notations. Let $\mathcal{V}$ denote the vocabulary and $\mathcal{V}^{*}$ the set of all possible textual inputs (Kleene closure of $\mathcal{V}$). Let $\mathcal{Q}\subset\mathcal{V}^{*}$ be the set of queries and $\mathcal{D}\subset\mathcal{V}^{*}$ the document database. In the reranking setting, a query $q\in\mathcal{Q}$ is associated with $k\in\mathbb{N}$ candidate documents $\left(d_{1},\cdots,d_{k}\right)\in\mathcal{D}^{k}$ coming from the preceding retrieval phase, some of them being relevant to the query and the others not. When available, we denote by $\mathbf{y}\in\left\{0,1\right\}^{k}$ the ground truth vector, such that $y_{j}=1$ if $d_{j}$ is relevant to $q$ and $y_{j}=0$ otherwise [^4] Note that $\mathbf{y}$ is not a one-hot vector as several documents can be relevant to a single query..
@@ 2.2 Abstention in Reranking
Requirements. In compliance with the set of constraints defined in the introduction (1), $c$ must not incur important extra computational costs, either during training or at inference time. The idea behind this requirement is that rerankers are for the most part already very large models, and we do not want to create an additional layer that makes them significantly heavier.
1. In this work, we therefore focus on the case in which the encoder parameters $\bm{\theta}$ are not learnable.
2. Furthermore, we constrain the choice of $c$ to the set of functions that only regard the relevance scores, i.e., $c\in\left\{u\circ g_{\bm{\theta}}\ |\ u:\mathbb{R}^{k}\rightarrow\mathbb{R}\right\}$.
@@ 2.3 Related Work
As far as our knowledge, no paper in the literature has specifically addressed the development of abstention mechanisms in the field of IR while respecting the black-box constraint. The closest initiatives encompass works proposing abstention strategies applied to general ranking cases, with the idea to allow a model to abstain at a certain cost during training by developing an approach based on partial rankings Cheng et al. (2010); Mao et al. (2023). As mentioned, however, these methods do not meet the no-training constraint.
More generally, abstention initiatives have been around for a long time in the classification setting Chow (1957), most of them being closely related to the notion of prediction confidence or uncertainty. Among the existing methods, three categories can be distinguished: learning-based, white-box, and black-box methods.
@@
Among learning-based methods, the most well-known remain Bayesian neural networks, which have a built-in capacity for uncertainty estimation Blundell et al. (2015). Similarly, aleatoric uncertainty estimation requires the use of a log-likelihood loss function Loquercio et al. (2020), thus greatly modifying the training process. Some approaches also include specific regularizers in the loss function Xin et al. (2021); Zeng et al. (2021), while others incorporate abstention as a class during training Rajpurkar et al. (2018), most of the time with a certain cost associated Bartlett and Wegkamp (2008); El-Yaniv and Wiener (2010); Cortes et al. (2016); Geifman and El-Yaniv (2019). In any case, these methods do not align with the no-training constraint.
@@
Most white-box methods, on the other hand, do not require specific training. Monte Carlo dropout is a popular approach involving randomly deactivating neurons of a model at inference time to estimate its epistemic uncertainty Houlsby et al. (2011); Gal et al. (2017); Smith and Gal (2018), while other approaches rely on a statistical analysis of the model’s features (hidden layers) Lee et al. (2018); Ren et al. (2021); Podolskiy et al. (2021); Haroush et al. (2021); Gomes et al. (2022). In both cases, white-box access to the model is required, which cannot be guaranteed (e.g., API services).
@@
Finally, black-box approaches assume only access to the output layer of the classifier, i.e., the soft probabilities. Some works estimate uncertainty by using the probability of the mode Hendrycks and Gimpel (2016); Lefebvre-Brossard et al. (2023); Hein et al. (2019); Hsu et al. (2020); Wang et al. (2023) while others utilize the entire logit distribution Liu et al. (2020). However, it is unclear how to apply these frameworks to the ranking scenario, where relevance scores are not soft probabilities.
@@ 3 Confidence Assessment for Document Reranking
Figure 1: Procedure diagram for black-box confidence estimation and abstention decision in a reranking setting. In the reference-free scenario, confidence function $u$ is a simple heuristic (e.g., maximum). In the reference-based scenario, $u$ is a light non-trivial function of the relevance scores (e.g., learned linear combination).
In this paper, as stated in Section 2.2, the confidence scorer we build only takes into account the vector of relevance scores. The intuition behind such an approach is that it is possible, given the distribution of these scores, to estimate the degree of confidence of our ranker on a given instance. In the next two subsections, we describe two methods to calculate this confidence score: the first being reference-free, and the second one requiring an access to a reference dataset.
@@ 3.1 Reference-Free Scenario
First, we present the reference-free approach, which can be divided into three stages (Figure 1).
1. Observation. We first observe the relevance scores $\mathbf{z}=g_{\bm{\theta}}(x)$ (Equation 2) for a given test instance $x=\left(q,d_{1},\cdots,d_{k}\right)\in\mathcal{Q}\times\mathcal{D}^{k}$.
2. Estimation. We then assess confidence using a simple heuristic function $u_{\text{base}}$, based solely on these relevance scores (e.g., maximum).
3. Thresholding. Finally, we compare the confidence score computed in step 2 to a threshold $\tau$. If $u_{\text{base}}(\mathbf{z})>\tau$, we make a prediction using ranker $r_{\bm{\theta}}$, otherwise, we abstain (Equation 5).
In this work, we evaluate a bunch of reference-free confidence functions, relying on simple statistics computed on the vector of relevance scores. We focus more particularly on three functions inspired from MSP Hendrycks and Gimpel (2016): (i) the maximum relevance score $u_{\text{max}}$, (ii) the standard deviation of the relevance scores $u_{\text{std}}$, and (iii) the difference between the first and second highest relevance scores $u_{\text{1-2}}$ Narayanan et al. (2012).
@@ 3.2 Data-Driven Scenario
An intuitive approach is to fit a simple supervised model whose objective is to predict ranking performance $m\left(g_{\bm{\theta}}(x),\mathbf{y}\right)$ given a vector of sorted relevance scores $s_{\bm{\theta}}(x)$. In this section, we develop a regression-based approach (see Figure 1).
0. Calibration. We use reference set $\mathcal{S}$ to derive $\mathcal{S}_{\bm{\theta}}$ (Equation 6) and then fit a regressor $h_{\text{reg}}:\mathbb{R}^{k}\rightarrow\mathbb{R}$ on $\mathcal{S}_{\bm{\theta}}$. Next, we derive $u_{\text{reg}}=h_{\text{reg}}\circ s$, the function that takes a vector of unsorted scores as input and returns the predicted ranking quality.
1. Observation. As in 3.1, we observe the relevance scores $\mathbf{z}=g_{\bm{\theta}}(x)$ for a given test instance $x$.
2. Estimation. We then assess confidence using the $u_{\text{reg}}$ function. Intuitively, the greater $u_{\text{reg}}(\mathbf{z})$, the more confident we are that $r_{\bm{\theta}}$ correctly ranks the documents with respect to the query.
3. Thresholding. As in 3.1, an abstention decision is finally made by comparing $u_{\text{reg}}(\mathbf{z})$ to $\tau$. In Section 5.3, we propose an approach to choose $\tau$.
In this work, we use a linear-regression-based confidence function $u_{\text{lin}}$ Fisher (1922). Formally, for a given unsorted vector of relevance scores $\mathbf{z}$,
@@
where $\beta_{0},\cdots,\beta_{k}\in\mathbb{R}$ are the coefficients fitted on $\mathcal{S}_{\bm{\theta}}$, with an $l_{2}$ regularization parameter $\lambda=0.1$ Hoerl and Kennard (1970) [^6] Scikit-learn implementation Pedregosa et al. (2011).. More data-driven confidence functions are explored in Appendix D.
@@ 4 Experimental Setup
In this section, we propose a novel experimental setup for assessing the performance of abstention mechanisms in a black-box reranking setting. We present the datasets and models that make up our benchmark, as well as the evaluation methods used.

### v5

@@ Abstract
Neural Information Retrieval (NIR) has significantly improved upon heuristic-based Information Retrieval (IR) systems. Yet, failures remain frequent, the models used often being unable to retrieve documents relevant to the user’s query. We address this challenge by proposing a lightweight abstention mechanism tailored for real-world constraints, with particular emphasis placed on the reranking phase. We introduce a protocol for evaluating abstention strategies in black-box scenarios (typically encountered when relying on API services), demonstrating their efficacy, and propose a simple yet effective data-driven mechanism. We provide open-source code for experiment replication and abstention implementation, fostering wider adoption and application in diverse contexts.
@@ 1 Introduction
Figure 1: Procedure diagram for black-box confidence estimation and abstention decision in a reranking setting. In the reference-free scenario, confidence function $u$ is a simple heuristic (e.g., maximum). In the data-driven scenario, $u$ is a light non-trivial function of the relevance scores (e.g., learned linear combination).
@@
In recent years, NIR has emerged as a promising approach to addressing the challenges of IR on various tasks (Guo et al., 2016; Mitra & Craswell, 2017; Zhao et al., 2022). Central to the NIR paradigm are the pivotal stages of retrieval and reranking (Robertson et al., 1995; Ni et al., 2021; Neelakantan et al., 2022), which collectively play a fundamental role in shaping the performance and outcomes of IR systems (Thakur et al., 2021). While the objective of the retrieval stage is to efficiently fetch candidate documents from a vast corpus based on a user’s query (frequentist (Ramos et al., 2003; Robertson et al., 2009) or bi-encoder-based approaches (Karpukhin et al., 2020)), reranking aims at reordering these retrieved documents in a slower albeit more effective way, using more sophisticated techniques (bi-encoder- or cross-encoder-based (Nogueira & Cho, 2019; Khattab & Zaharia, 2020)). [^1] The most common approach is to have a quick retriever that scales well and a more computationally intensive reranker that reorders the top retrieved contexts.
@@
However, despite advancements in NIR techniques, retrieval and reranking processes frequently fail, due for instance to the quality of the models being used, to ambiguous user queries or insufficient relevant documents in the database. These can have unsuitable consequences, particularly in contexts like Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), where retrieved information is directly used in downstream tasks, thus undermining the overall system reliability (Yoran et al., 2023; Wang et al., 2023; Baek et al., 2023).
@@
Recognizing the critical need to address these challenges, the implementation of abstention mechanisms emerges as a significant promise within the IR landscape. This technique offers a pragmatic approach by abstaining from delivering results when the model shows signs of uncertainty, thus ensuring users are not misled by potentially erroneous information. Although the subject of abstention in machine learning has long been of interest (Chow, 1957; El-Yaniv & Wiener, 2010; Geifman & El-Yaniv, 2017; 2019), most research has focused on the classification setting, leaving lack of significant initiatives in IR. A few related works propose computationally intensive learning strategies (Cheng et al., 2010; Mao et al., 2023) that are difficult to implement in the case of NIR, where models can have up to billions of parameters.
@@
Scope. In this paper, we explore lightweight abstention mechanisms that adhere to realistic industrial constraints, namely (i) black-box access-only to query-documents relevance scores, which is typically the case when relying on API services, (ii) marginal computational and latency overhead, (iii) configurable abstention rates depending on the application. In line with industrial applications of IR systems for RAG or search engines, our method focuses on assessing the reranking quality of the top retrieved documents.
Contributions. Our key contributions are:
  1. 1.
@@
    We devise a detailed protocol for evaluating abstention strategies in a black-box reranking setting and demonstrate their relevance.
  2. 2.
@@
    We introduce a simple yet effective reference-based abstention mechanism, outperforming reference-free heuristics at zero incurred cost. We perform ablations and analyses on this method and are confident about its potential in practical settings.
  3. 3.
@@
    We release a code package [^2] https://github.com/artefactory/abstention-reranker, under MIT license. and artifacts [^3] https://huggingface.co/collections/artefactory/abstention-reranking. to enable full replication of our experiments and the implementation of plug-and-play abstention mechanisms for any use case.
@@ 2.1 Notations
General notations. Let $\mathcal{V}$ denote the vocabulary and $\mathcal{V}^{*}$ the set of all possible textual inputs (Kleene closure of $\mathcal{V}$). Let $\mathcal{Q}\subset\mathcal{V}^{*}$ be the set of queries and $\mathcal{D}\subset\mathcal{V}^{*}$ the document database. In the reranking setting, a query $q\in\mathcal{Q}$ is associated with $k\in\mathbb{N}$ candidate documents $\left(d_{1},\cdots,d_{k}\right)\in\mathcal{D}^{k}$ coming from the preceding retrieval phase, some of them being relevant to the query and the others not. When available, we denote by $\mathbf{y}\in\left\{0,1\right\}^{k}$ the ground truth vector, such that $y_{j}=1$ if $d_{j}$ is relevant to $q$ and $y_{j}=0$ otherwise. [^4] Note that $\mathbf{y}$ is not a one-hot vector as several documents can be relevant to a single query.
@@ 2.2 Abstention in Reranking
Requirements. In compliance with the set of constraints defined in the introduction (Section 1), $c$ must be usable in a fully black-box setting and not incur significant extra computational costs, either during training or at inference time. Formally, we set the two following requirements:
  1. 1.
    We focus on the case in which the encoder parameters $\bm{\theta}$ are not learnable.
  2. 2.
    Furthermore, $c$ must only take relevance scores as input, i.e., $c\in\left\{u\circ g_{\bm{\theta}}\ |\ u:\mathbb{R}^{k}\rightarrow\mathbb{R}\right\}$.
@@ 2.3 Related Work
Abstention, also known as selective prediction, has been around for a long time, originally for classification purposes (Chow, 1957; El-Yaniv & Wiener, 2010), and has often been associated with topics such as confidence estimation, out-of-domain (OOD) detection (Schölkopf et al., 1999; Liang et al., 2017; Darrin et al., 2022; 2024) and prediction error detection (Hendrycks & Gimpel, 2016). These fields collectively encompass three main approaches: learning-based methods, which require a dedicated training process; white-box methods, which rely on full access to the model’s internal features; and black-box methods, which focus solely on the model’s output.
@@
Among learning-based methods, the most classic initiatives include algorithms such as SVMs, nearest neighbors, boosting (Hellman, 1970; Fumera & Roli, 2002; Wiener & El-Yaniv, 2015; Cortes et al., 2016) or Bayesian methods (Blundell et al., 2015) such as Markov Chain Monte Carlo (Geyer, 1992) and Variational Inference (Hinton & Van Camp, 1993; Graves, 2011), that include uncertainty estimation as a whole part of the training process. Similarly, aleatoric uncertainty estimation requires the use of a log-likelihood loss function (Loquercio et al., 2020), thus greatly modifying the training process. Some approaches also include specific regularizers in the loss function (Xin et al., 2021; Zeng et al., 2021), while others incorporate abstention as a class during training (Rajpurkar et al., 2018), most of the time with a certain cost associated (Bartlett & Wegkamp, 2008; El-Yaniv & Wiener, 2010; Cortes et al., 2016; Geifman & El-Yaniv, 2019). It comes obvious that none of these methods align with the no-training requirement set in Section 2.2.
@@
Most white-box methods, on the other hand, do not require specific training. When considering selective prediction in neural networks, a straightforward and effective technique is to set a threshold over a confidence score derived from a pre-trained network, which was already optimized to predict all points (Cordella et al., 1995; De Stefano et al., 2000; Wiener & El-Yaniv, 2012). Monte Carlo dropout is another popular approach involving randomly deactivating neurons of a model at inference time to estimate its epistemic uncertainty (Houlsby et al., 2011; Gal et al., 2017; Smith & Gal, 2018). Others rely on a statistical analysis of the model’s features (hidden layers) (Lee et al., 2018; Ren et al., 2021; Podolskiy et al., 2021; Haroush et al., 2021; Gomes et al., 2022) while ensemble-based methods (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017; Geifman & El-Yaniv, 2019) extend this idea and estimate confidence based on statistics of the ensemble model’s features. The latter approaches might significantly increase inference time, which is not in line with our set of constraints (Section 2.2). Moreover, white-box access to the model cannot always be guaranteed in practical situations (e.g., API services).
@@
Black-box approaches assume only access to the output layer of the classifier, i.e., the soft probabilities or logits. Some works estimate uncertainty by using the probability of the mode (Hendrycks & Gimpel, 2016; Lefebvre-Brossard et al., 2023; Hein et al., 2019; Hsu et al., 2020; Wang et al., 2023) while others utilize the entire logit distribution (Liu et al., 2020), the most widely known and straightforward confidence calibration technique remaining temperature scaling (Platt et al., 1999). However, it is unclear how to apply these frameworks to the ranking scenario, where relevance scores are neither soft probabilities nor logits but unscaled scalars.
Selective prediction and confidence estimation in NLP have seen targeted applications but have room for expansion across broader scenarios. Dong et al. (2018) describe a model developed to estimate confidence specifically for semantic parsing, while Kamath et al. (2020) delve into selective prediction for out-of-distribution (OOD) question answering, allowing the model to abstain from answering questions deemed too difficult or outside the training distribution. More recently, Xin et al. (2021) have applied selective prediction through loss regularization in various classification settings. Additionally, datasets such as SQuAD2.0 (Rajpurkar et al., 2018) illustrate the challenge of unanswerable questions. In the IR field specifically, strategies like those proposed by Cheng et al. (2010) and Mao et al. (2023) allow a model to abstain at a certain cost during training by assessing the confidence level on partial ranks. Here again, such methods do not fit the black-box constraint mentioned in Section 2.2.
@@ 3 Confidence Assessment for Document Reranking
In this paper, as stated in Section 2.2, the confidence scorer we build only takes into account the list of relevance scores. Our intuition is that, by analyzing the distribution of the ranker’s scores, it should be possible to estimate its confidence on a given instance. In the next two subsections, we describe two methods for calculating the confidence score: the first one is reference-free, meaning it does not require any external data, while the second one relies on access to reference data for calibration purposes.
@@ 3.1 Reference-Free Scenario
First, we present the reference-free approach, which can be divided into three stages (Figure 1):
  1. 1.
    Observation. We first observe the relevance scores $\mathbf{z}=g_{\bm{\theta}}(x)$ (Equation 2) for a given test instance $x=\left(q,d_{1},\cdots,d_{k}\right)\in\mathcal{Q}\times\mathcal{D}^{k}$.
  2. 2.
    Estimation. We then assess confidence using a simple heuristic function $u_{\text{base}}$, based solely on these relevance scores (e.g., maximum).
  3. 3.
    Thresholding. Finally, we compare the confidence score computed in step 2 to a threshold $\tau$. If $u_{\text{base}}(\mathbf{z})>\tau$, we make a prediction using ranker $r_{\bm{\theta}}$, otherwise, we abstain (Equation 5).
In this work, we evaluate a bunch of reference-free confidence functions, relying on simple statistics computed on the relevance scores. We focus more particularly on three functions inspired from MSP (Hendrycks & Gimpel, 2016): (i) the maximum relevance score $u_{\text{max}}$, (ii) the standard deviation $u_{\text{std}}$, and (iii) the difference between the first and second highest relevance scores $u_{\text{1-2}}$ (Narayanan et al., 2012).
@@ 3.2 Data-Driven Scenario
An intuitive approach is to fit a simple supervised model to predict ranking performance $m\left(g_{\bm{\theta}}(x),\mathbf{y}\right)$ given a vector of sorted relevance scores $s_{\bm{\theta}}(x)$. In this section, we develop a regression-based approach (Figure 1):
  1. 0.
    Calibration. We use reference set $\mathcal{S}$ to derive $\mathcal{S}_{\bm{\theta}}$ (Equation 6) and then fit a regressor $h_{\text{reg}}:\mathbb{R}^{k}\rightarrow\mathbb{R}$ on $\mathcal{S}_{\bm{\theta}}$. Next, we derive $u_{\text{reg}}=h_{\text{reg}}\circ s$, the function that takes a vector of unsorted scores as input and returns the predicted ranking quality.
  2. 1.
    Observation. As in Section 3.1, we observe the relevance scores $\mathbf{z}=g_{\bm{\theta}}(x)$ for a given test instance $x$.
  3. 2.
    Estimation. We then assess confidence using the $u_{\text{reg}}$ function. Intuitively, the greater $u_{\text{reg}}(\mathbf{z})$, the more confident we are that $r_{\bm{\theta}}$ correctly ranks the documents with respect to the query.
  4. 3.
    Thresholding. As in Section 3.1, an abstention decision is finally made by comparing $u_{\text{reg}}(\mathbf{z})$ to $\tau$. In Section 5.3, we propose an approach to choose $\tau$.
In this work, we use a linear-regression-based confidence function $u_{\text{lin}}$ (Fisher, 1922). Formally, for a given unsorted vector of relevance scores $\mathbf{z}$,
@@
where $\beta_{0},\cdots,\beta_{k}\in\mathbb{R}$ are the coefficients fitted on $\mathcal{S}_{\bm{\theta}}$, with an $l_{2}$ regularization parameter $\lambda=0.1$ (Hoerl & Kennard, 1970). [^6] Scikit-learn implementation (Pedregosa et al., 2011). More data-driven confidence functions are explored in Appendix D.
@@ 4 Experimental Setup
In this section, we propose a novel experimental setup for assessing performance of abstention mechanisms in a black-box reranking setting. We present both our benchmark and evaluation metrics.